Bobby is 6 years older than Daniel. If you take 5 times bobby's age and add 2 times Daniel's age the result is 247. How old are they?
Answer by addingup(3677)

B = D+6 We'll use this value for B next:
5B+2D = 247 Now substitute for B:
5(D+6)+2D = 247
5D+30+2D = 247
7D = 217 Divide both sides by 7:
D = 31 and B:
31+6 = 37
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
"If you take 5 times bobby's age and add 2 times Daniel's age the result is 247"
5(37)+2(31)= 247
185+62 = 247 We have the correct answer.